### v1

@@ 3.3 Transfer Operator with Circle rotation
with orthogonal $Q_{\Phi},Q_{\Psi}\in\mathbb{R}^{n\times n}$ and upper-triangular $R_{\Phi},R_{\Psi}\in\mathbb{R}^{n\times N}$, denote by $Q_{\Phi}^{(r)}$ and $Q_{\Psi}^{(r)}$ the first $r$ columns of $Q_{\Phi}$ and $Q_{\Psi}$. If $0\leq\theta_{1}\leq\cdots\leq\theta_{r}$ are the principal angles between the subspaces spanned by $\Phi$ and $\Psi$, then
@@
From these angles we report three standard subspace‑distance metrics:
@@
Table 4: Subspace‐distance measures between $\operatorname{span}\{\phi_{j}\}_{j=1}^{N}$ and $\operatorname{span}\{\mathcal{A}\phi_{j}\}_{j=1}^{N}$
Metric $\beta_{3}=0$ $\beta_{3}=0.1$ $\beta_{3}=0.2$ $\beta_{3}=0.6$
Geodesic ($d_{\max}$) $4.88\times 10^{-4}$ $4.88\times 10^{-4}$ $5.98\times 10^{-4}$ $5.98\times 10^{-4}$
Chordal ($d_{\text{ch}}$) $1.38\times 10^{-3}$ $1.38\times 10^{-3}$ $1.54\times 10^{-3}$ $1.29\times 10^{-3}$
Projection ($d_{\text{pr}}$) $4.88\times 10^{-4}$ $4.88\times 10^{-4}$ $5.98\times 10^{-4}$ $5.98\times 10^{-4}$
Smaller values of the principal angles $\theta_{k}$ indicate that the two subspaces are closer, and hence the learned subspace approaches invariance under $A$. Across all sparsity levels $\beta_{3}$, the reported distances remain on the order of $10^{-3}$ or less, providing strong evidence of a near-invariant learned subspace.
3.4 Transfer Operator with Perturbed Arnold’s Cat Map
@@
with $\delta=0.01$. With this perturbation, $T$ remains Anosov and therefore has an SRB measure $\mu$ [5, Theorem 4.12]. The spectral theory for Anosov diffeomorphisms is challenging and relies on carefully constructed anisotropic Banach spaces (e.g. [13]) that are adapted to the unstable and stable foliations. In this example, we let $\mathcal{L}f=f\circ T^{-1}/|\det DT\circ T^{-1}|$ be the Perron–Frobenius operator, and consider $\mathcal{L}:L^{2}(X,m)\to L^{2}(X,m)$ where $m$ is Lebesgue measure on the flat torus. Note the additional determinant term in the definition of $\mathcal{L}$; the constant function is no longer the leading eigenfunction of $\mathcal{L}$. In fact, the object $\mu$ satisfying $\mathcal{L}\mu=\mu$ is in general a distribution and need not lie in $L^{2}(\mathbb{T}^{2},m)$. The probability measure (or distribution) $\mu$ describes the long-term distribution of infinitely long trajectories of $T$ for Lebesgue almost all initial conditions $x\in\mathbb{T}^{2}$. This distribution of infinitely long trajectories is accessible through the fixed point of $\mathcal{L}$; this is one of the many advantages of studying transfer operators.
To eliminate the discontinuities that arise at the edges of the unit square, we embed $\mathbb{T}^{2}$ into $\mathbb{R}^{4}$ by mapping each $(x,y)$ to its pair of complex phases $(\cos 2\pi x,\allowbreak\sin 2\pi x,\allowbreak\cos 2\pi y,\allowbreak\sin 2\pi y)$. We first evaluate our model’s ability to approximate the Perron–Frobenius operator on an unseen observable. Throughout this entire example no sparsity penalty is applied ($\beta_{3}=0$). Fig. 8 juxtaposes the input observable $g$, the ground-truth image, and our network’s prediction. The accuracy is quantified by the relative L2 error of $1.17\times 10^{-2}$ on the full test set.
Figure 8: Input–output approximation for the Perron–Frobenius operator of a perturbed Arnold’s cat map. The three panels show the unseen input observable $g$, the corresponding ground-truth response, and the network prediction. The mean $L^{2}$ relative error over the full test set is $1.17\times 10^{-2}$.
Next, we analyse the learned set of basis functions that span the invariant subspace $V=\operatorname{span}\{\phi_{j}\}_{j=1}^{N}$. Fig. 9 shows representative basis functions obtained without sparsity penalty. Even in the absence of an explicit localisation term, the basis functions obtained through the proposed network show unmistakable alignment with contracting and expanding directions.
@@
The Gram matrix of the learned basis functions is again defined by
$G_{kj}=\langle\phi_{k},\phi_{j}\rangle,\qquad k,j=1,\dots,N,$
and is illustrated in Fig. 10. Its nearly diagonal form shows that our training scheme implicitly promotes approximate orthogonality, even in the absence of an explicit sparsity term.
Figure 10: Gram matrix of the learned basis functions (after $\ell^{2}$ normalisation). The (approximately) diagonal structure shows that the network has produced a nearly orthogonal collection, despite the absence of an explicit sparsity constraint.
We now examine the spectral properties of the learned Perron–Frobenius approximator. Fig. 11 displays the eigenvalues; the leading eigenvalue at $\lambda=1$ confirms the presence of an invariant measure consistent with the SRB measure associated with this map.
@@
The remaining eigenvalues in Fig. 11 are restricted to a disk away from the unit circle, in line with expectations if one used a correct anisotropic setup [13]. While we work in $L^{2}(X,m)$, we train our estimate of $\mathcal{L}$ on smooth functions, and it is likely that this restriction to smooth functions produces the spectral gap we see in Fig. 11.
Using the same Galerkin projection technique as in the previous example, we reconstruct the leading eigenfunction that represents the SRB measure and, for comparison, include the same leading eigenfunction reconstructed from a Fourier basis of the same dimension — built from products of sines and cosines described in Section 3.1; see Fig. 12.
Figure 12: Ground-truth SRB measure (left), its reconstruction using the SABON learned basis (middle), and its reconstruction using a Fourier basis of the same dimension (right).
@@
To benchmark the learned basis against the traditional Fourier basis used to construct the trigonometric polynomials in Section 3.1, we measure the relative $L^{2}$-projection error of the ground-truth SRB measure,
$\varepsilon_{\mathcal{B}}^{\mathrm{rel}}:=\frac{\|\rho_{\mathrm{SRB}}-P_{\mathcal{B}}\rho_{\mathrm{SRB}}\|_{2}}{\|\rho_{\mathrm{SRB}}\|_{2}},\quad P_{\mathcal{B}}:\;L^{2}(X)\rightarrow\operatorname{span}(\mathcal{B})\text{ is the orthogonal projection.}$.
We compute this error for (i) the learned basis and (ii) a Fourier basis of comparable dimension. The learned basis achieves the smaller error (see Table 5).
Table 5: Relative orthogonal projection error $\varepsilon_{\mathcal{B}}^{\mathrm{rel}}=\frac{\|\rho_{\mathrm{SRB}}-P_{\mathcal{B}}\rho_{\mathrm{SRB}}\|_{2}}{\|\rho_{\mathrm{SRB}}\|_{2}}$. Lower values indicate a more accurate representation of the invariant measure.
Basis Relative projection error
SABON (225 functions) $7.248\times 10^{-2}$
Fourier (225 functions) $1.318\times 10^{-1}$
Finally, we quantify and compare the invariance of the subspaces generated by the learned basis and a classical Fourier basis of the same size under the action of the Perron–-Frobenius operator. We compute principal angles between each subspace and its image under the operator, summarising these in Table 6. The significantly smaller distances obtained for our learned subspace indicate superior invariance compared with the traditional Fourier basis.
@@
Geodesic ($d_{\max}$) $0.088$ $16.642$
Chordal ($d_{\text{ch}}$) $0.088$ $10.615$
Projection ($d_{\text{pr}}$) $0.087$ $1.000$
@@
Figure 13: Projection distance for the principal angles between each subspace and its image under the Perron–Frobenius operator for the learned basis and the Fourier basis of equal dimension. Smaller distances correspond to greater invariance.
Why does the SABON basis have far superior invariance when compared to a Fourier basis? A Fourier element—-whether the complex exponential $e^{2\pi ik\cdot x}$ or the purely real sine–cosine products introduced in Section 3.1—-is aligned with coordinate axes. The contracting dynamics of $T$ means that the frequency of a Fourier mode will be increased in stable directions of the dynamics under application of $\mathcal{L}$. Thus those modes in the Fourier basis with higher frequencies may be pushed far out of the original span of the Fourier modes, and made near-orthogonal to the Fourier basis. Indeed, Fig. 13 indicates that this occurs to more than half of the Fourier modes.
The SABON basis, on the other hand, is learned jointly with the dynamics, so the network discovers functions that organise along the stable and unstable directions. The resulting modes resemble elongated ridges: long in the expanding direction and sharply confined in the contracting one. When such a ridge is pushed forward by $\mathcal{L}$ it tends to become another ridge, aiding the invariance of the SABON subspace. This anisotropic feature, absent from the isotropic Fourier family, accounts for the much smaller distances reported for SABON in Table 6 and illustrated in Fig. 13.
@@ 4 Conclusions
We have introduced the Single Autoencoder Basis Operator Network (SABON)—a neural framework that jointly learns
@@
SABON cleanly achieves this by combining an encoder–decoder structure with a linear latent map $\mathcal{G}$ and a composite loss that balances reconstruction accuracy, operator approximation accuracy, and sparsity. We proved a universal approximation theorem in the spirit of Hua [14], but specialised to a single neural basis that targets transfer and Koopman operators.
@@
    Circle rotation. The SABON basis provided accurate reconstruction of the analytic eigenpairs and recovered the rotation angle $\alpha$ directly from the leading eigenvalues, while the associated eigenfunctions coincide with the theoretical ${e^{\,ik\theta}}$ modes. This example provided evidence that the learned basis accurately captured the main dynamical features of the system.
@@
    Perturbed Arnold cat map. For a nonlinear uniformly hyperbolic map SABON’s learned basis outperformed a traditional Fourier basis, by adapting to the geometry inherent in the dynamics. This is particularly surprising given the well-known theoretical approximation strengths of Fourier modes. The network discovered basis functions largely aligned with the contracting and expanding directions, leading to more accurate estimates of the physical invariant measure of the dynamics and a highly invariant approximation subspace.
Future directions
Promising paths for further research include:
  • learning the operator spectrum directly in latent space, bypassing external Galerkin projections;
  • investigating additional examples where isotropic bases fail and data-adaptive bases excel;
  • extending SABON to high-dimensional or partially observed systems to assess scalability and sample efficiency;
  • deriving quantitative rates that relate network width, depth and sparsity to the operator approximation error.

### v2

@@ 3.3 Transfer operator for a circle rotation
with orthogonal $Q_{\Phi},Q_{\Psi}\in\mathbb{R}^{n\times n}$ and upper-triangular $R_{\Phi},R_{\Psi}\in\mathbb{R}^{n\times N}$. Denote by $Q_{\Phi}^{(r)}$ and $Q_{\Psi}^{(r)}$ the first $r$ columns of $Q_{\Phi}$ and $Q_{\Psi}$. If $0\leq\theta_{1}\leq\cdots\leq\theta_{r}$ are the principal angles between the subspaces spanned by $\Phi$ and $\Psi$, then
@@
To avoid numerical issues associated with the classical cosine-based computation for small angles and the sine-based approach for angles near $\frac{\pi}{2}$, we employed the combined technique [20] that uses the cosine-based method initially and switches to the sine-based approach for principal angles smaller than $\frac{\pi}{4}$. Note that the computation of principal angles is extremely sensitive to numerical precision. Therefore, we use higher precision floating-point arithmetic for these operations in both of our examples to ensure accurate results.
From these angles $\theta=(\theta_{1},\ldots,\theta_{N})$, we report three standard subspace‑distance metrics:
@@
Table 4: Subspace‐distance measures between $\operatorname{span}\{\phi_{j}\}_{j=1}^{N}$ and $\operatorname{span}\{\mathcal{L}(\phi_{j})\}_{j=1}^{N}$
Metric $\beta_{2}=0$ $\beta_{2}=0.1$ $\beta_{2}=0.2$ $\beta_{2}=0.6$
Geodesic ($d_{\max}$) $3.983\times 10^{-2}$ $3.621\times 10^{-2}$ $3.869\times 10^{-2}$ $4.008\times 10^{-2}$
Chordal ($d_{\text{ch}}$) $3.982\times 10^{-2}$ $3.621\times 10^{-2}$ $3.868\times 10^{-2}$ $4.008\times 10^{-2}$
Projection ($d_{\text{pr}}$) $1.760\times 10^{-2}$ $1.906\times 10^{-2}$ $2.146\times 10^{-2}$ $1.732\times 10^{-2}$
Smaller values of the principal angles $\theta_{k}$ indicate that the two subspaces are closer, and hence the learned subspace approaches invariance under $\mathcal{L}$. Across all sparsity levels $\beta_{2}$, the reported distances remain on the order of $10^{-2}$, providing strong evidence of a learned subspace that is nearly invariant under the true operator $\mathcal{L}$.
We close this example by discussing the choice of $N=19$ basis functions. Recall we trained $\hat{\mathcal{L}}$ using trigonometric polynomials up to order 9 (equivalently linear combinations of real Fourier modes up to order 9). As discussed above, each two-dimensional real subspace spanned by $\theta\mapsto\sin(k\theta)$ and $\theta\mapsto\cos(k\theta)$ is an eigenspace with eigenpair $\lambda_{\pm k}$. Thus, in this special situation of circle rotation dynamics, the input and output data lie in the same 19-dimensional (19=2×9+119=2$\times$ 9+119 = 2 × 9 + 1) subspace. Increasing the learned basis cardinality $N$ beyond 19 will not impact the prediction error, but the principal angles beyond the first 19 will increase because these additional dimensions will effectively be unable to be controlled with $9^{\rm th}$-order input/output data.
3.4 Transfer operator for a nonlinearly perturbed Arnold’s cat map
@@
with $\delta=0.01$. With this perturbation, $T$ remains Anosov and therefore has an SRB measure $\mu$ [5, Theorem 4.12]. The spectral theory for Anosov diffeomorphisms is challenging and relies on carefully constructed anisotropic Banach spaces (e.g. [15]) that are adapted to the unstable and stable foliations. In this example, we let $\mathcal{L}f=f\circ T^{-1}/|\det DT\circ T^{-1}|$ be the Perron–Frobenius operator, and consider $\mathcal{L}:L^{2}(X,m)\to L^{2}(X,m)$ where $m$ is Lebesgue measure on the flat torus. Note the additional determinant term in the definition of $\mathcal{L}$; the constant function is no longer the leading eigenfunction of $\mathcal{L}$. In fact, the object $\mu$ satisfying $\mathcal{L}\mu=\mu$ is in general a distribution and need not lie in $L^{2}(\mathbb{T}^{2},m)$. The fixed distribution $\mu$ may be identified with the SRB measure and describes the long-term distribution of infinitely long trajectories of $T$ for Lebesgue almost all initial conditions $x\in\mathbb{T}^{2}$. This distribution of infinitely long trajectories is accessible through the fixed point of $\mathcal{L}$; this is one of the many advantages of studying transfer operators.
[…]
Figure 8: Input and output for the Perron–Frobenius operator of a perturbed Arnold’s cat map. Left to right, the three panels show (i) an unseen input observable $g$, (ii) the corresponding ground-truth output from the true transfer operator, (iii) and the corresponding predicted output from our learned transfer operator. The mean $L^{2}$ relative error over the full test set is $1.225\times 10^{-2}$.
Next, we analyse the learned set of basis functions $\{\phi_{j}\}_{j=1}^{N}$ that span the invariant subspace $V$. Fig. 9 shows representative basis functions obtained without sparsity penalty. The basis functions obtained through the proposed network show unmistakable alignment with contracting and expanding directions of the nonlinear dynamics generated by $T$.
@@
The Gram matrix of the $L^{2}$-normalised learned basis functions defined in (7) is illustrated in Fig. 10. Its nearly diagonal form shows that our training scheme implicitly promotes approximate orthogonality, even without an explicit sparsity term.
Figure 10: Gram matrix of the learned basis functions (after $L^{2}$ normalisation). The diagonal structure shows that the network has produced a nearly orthogonal collection, despite the absence of an explicit sparsity constraint.
We now examine the spectral properties of the learned Perron–Frobenius approximator. Fig. 11 displays the eigenvalues; the leading eigenvalue at $\lambda=1$ confirms the presence of an invariant measure as a fixed point of the Perron–Frobenius operator.
@@
The remaining eigenvalues in Fig. 11 are restricted to a disk away from the unit circle, in line with expectations if one used a suitable anisotropic Banach space for the domain of $\mathcal{L}$ [15]. While we work in $L^{2}(X,m)$, we train our estimate of $\mathcal{L}$ on smooth functions, and it is likely that this restriction to smooth functions produces the spectral gap we see in Fig. 11.
Solving the eigenequation (8) similarly to the previous example, we reconstruct the leading eigenfunction (corresponding to $\lambda=1$), which represents the SRB measure; see Fig. 12.
Figure 12: Densities of ground-truth SRB measure (left), its reconstruction using the SABON learned basis (middle), and its reconstruction using a Fourier basis of the same dimension (right). We consider the densities as vectors on the grid and make their $\ell_{1}-$norms equal to 1.
@@
To quantitatively benchmark the learned basis against the traditional Fourier basis we measure both the relative $L^{2}$ error and Wasserstein $W_{1}$ distance of the orthogonally projected ground-truth SRB measure onto each basis. Letting $\mathcal{B}$ denote either of the two bases, denote the orthogonal projector onto $\mathcal{B}$ by $P_{\mathcal{B}}:L^{2}(X)\to\operatorname{span}(\mathcal{B})$. Define the relative $L^{2}$ projection error
$L^{2,\mathrm{rel}}_{\mathcal{B}}\bigl{(}\rho_{\mathrm{SRB}},P_{\mathcal{B}}(\rho_{\mathrm{SRB}})\bigr{)}\;\coloneq\;\frac{\|\rho_{\mathrm{SRB}}-P_{\mathcal{B}}(\rho_{\mathrm{SRB}})\|_{2}}{\|\rho_{\mathrm{SRB}}\|_{2}}.$
An SRB measure is typically a distribution and not a function, and therefore it has no density with respect to Lebesgue measure. This makes comparisons in the $L^{2}$ norm less meaningful (the $L^{2}$ topology is too strong). More appropriate is the 111-Wasserstein metric on probability measures. The 1-Wasserstein distance is defined as
$W_{1}(\rho,\eta)\;\coloneq\;\inf_{\pi\in\Pi(\rho,\eta)}\int_{X\times X}d(x,y)\,\mathrm{d}\pi(x,y),$
where $d(x,y)$ is the metric distance between $x$ and $y$ on the phase space $X$ and $\Pi(\rho,\eta)$ denotes all probability measures $\pi$ on $X\times X$ whose first marginal is $\rho$ and whose second marginal is $\eta$. After discretising $\rho$ and $\eta$ onto the uniform grid $\{x_{i}\}_{i=1,\dots,n}$, the infimum above can be written as the linear program
$\min_{\pi\geq 0}\;\sum_{i,j=1}^{n}\pi_{ij}C_{ij}\quad\text{s.t.}\quad\sum_{j=1}^{n}\pi_{ij}=\rho_{i},\;\sum_{i=1}^{n}\pi_{ij}=\eta_{j},$
where $C_{ij}$ is the distance between $x_{i}$ and $x_{j}$ in $X$ for $i,j=1,\ldots,n$, $\pi_{ij}$ is the mass flow between $x_{i}$ and $x_{j}$, and $0\leq\rho_{i},\eta_{i}\leq 1$ are the discrete masses assigned by $\rho,\eta$ (respectively) to the point $x_{i}\in X$ for each $i=1,\ldots,n$.
As in the relative $L^{2}$ norm calculation, we compute $W_{1}\bigl{(}\rho_{\mathrm{SRB}},P_{\mathcal{B}}(\rho_{\mathrm{SRB}})\bigr{)}$ when (i) $\mathcal{B}$ is the learned basis and (ii) $\mathcal{B}$ is a Fourier basis of comparable cardinality. Table 5 shows that in both $L^{2}$ and Wasserstein distances, the learned SABON basis achieves the smallest error, which implies that the learned basis is able to better capture the detailed structure in the “ground truth” SRB distribution.
Table 5: Approximation error of the ground-truth SRB measure under orthogonal projection onto the two candidate bases. Lower values denote a more accurate representation.
Basis $L^{2,\mathrm{rel}}_{\mathcal{B}}\bigl{(}\rho_{\mathrm{SRB}},P_{\mathcal{B}}(\rho_{\mathrm{SRB}})\bigr{)}$ $W_{1}\bigl{(}\rho_{\mathrm{SRB}},P_{\mathcal{B}}(\rho_{\mathrm{SRB}})\bigr{)}$
SABON (225 functions) $7.479\times 10^{-2}$ $7.886\times 10^{-4}$
Fourier (225 functions) $1.318\times 10^{-1}$ $1.717\times 10^{-3}$
Finally, we quantify and compare the invariance of the subspaces generated by the learned basis and a classical Fourier basis of the same size under the action of the Perron–Frobenius operator. We compute principal angles between each subspace and its image under the operator, summarising these in Table 6. The significantly smaller distances obtained for our learned subspace indicate superior invariance compared with the traditional Fourier basis.
@@
Geodesic ($d_{\max}$) 0.0370.0370.037 16.64216.64216.642
Chordal ($d_{\text{ch}}$) 0.0370.0370.037 10.61510.61510.615
Projection ($d_{\text{pr}}$) 0.0260.0260.026 1.0001.0001.000
@@
Figure 13: Projection distance for the principal angles between each subspace and its image under the Perron–Frobenius operator for the learned basis and the Fourier basis of equal cardinality. Smaller distances correspond to greater invariance.
Why does the SABON basis have far superior invariance when compared to a Fourier basis? A Fourier element — whether the complex exponential $e^{2\pi ik\cdot x}$ or the purely real sine–cosine products introduced in Section 3.1 — is aligned with coordinate axes. The contracting dynamics of $T$ means that the frequency of a Fourier mode will be increased in stable directions of the dynamics under application of $\mathcal{L}$. Thus, those modes in the Fourier basis with higher frequencies may be pushed far out of the original span of the Fourier modes, and made near-orthogonal to the Fourier basis. Indeed, Fig. 13 indicates that this occurs to more than half of the Fourier modes.
The SABON basis, on the other hand, is learned jointly with the dynamics, so the network discovers functions that organise along the stable and unstable directions. The resulting modes resemble elongated ridges: smooth in the expanding direction and more sharply confined in the contracting one. When such a ridge is pushed forward by $\mathcal{L}$ it tends to become another ridge, consistent with the invariance of the SABON subspace. This anisotropic feature, absent from the isotropic Fourier family, accounts for the much smaller distances reported for SABON in Table 6 and illustrated in Fig. 13.
@@ 4 Conclusions
We introduced the Single Autoencoder Basis Operator Network (SABON)—a neural framework that jointly learns
@@
SABON cleanly achieves this by combining an encoding–reconstruction structure with a linear latent map $\mathcal{G}$ and a composite loss that balances operator approximation accuracy and sparsity. We proved a universal approximation theorem in the spirit of Hua [16], specialised to a single neural basis that targets transfer and Koopman operators.
@@
    Circle rotation. The SABON-learned basis accurately reconstructed the analytic eigenpairs and recovered the rotation angle $\alpha$ directly from the leading eigenvalues, while the associated eigenfunctions coincided with the theoretical ${e^{\,ik\theta}}$ modes. Thus, the learned basis accurately captured the main dynamical properties.
@@
    Perturbed Arnold’s cat map. SABON’s learned basis outperformed a traditional Fourier basis for a nonlinear uniformly hyperbolic map by adapting to the geometry inherent in the dynamics. This is particularly surprising given the well-known theoretical approximation strengths of Fourier modes. The network discovered basis functions largely aligned with the contracting and expanding directions, leading to more accurate estimates of the physical invariant measure of the dynamics and a highly invariant approximation subspace.
SABON opens up new possibilities for constructing compact and accurate transfer and Koopman operator representations, particularly in difficult situations where isotropic bases fail and data-adaptive bases may excel. Moreover, SABON is in principle extendable to high-dimensional or partially observed systems. Future theoretical challenges include deriving quantitative rates that relate network width, depth, and sparsity to the operator approximation error.
Data availability
All code and data supporting our findings is publicly available in our repository at https://github.com/kevinkuhl/SABON.